# On the approximation of finite-time Lyapunov exponents for the stochastic Burgers equation

### v1

@@ Abstract
We analyze stochastic partial differential equations (SPDEs) with quadratic nonlinearities close to a change of stability. To this aim we compute finite-time Lyapunov exponents (FTLEs), observing a change of sign based on the interplay between the distance towards the bifurcation and the noise intensity. A technical challenge is to provide a suitable control of the nonlinear terms coupling the dominant and stable modes of the SPDE and of the corresponding linearization. In order to illustrate our results we apply them to the stochastic Burgers equation.
Keywords: finite-time Lyapunov exponents, amplitude equations, Burgers equation.
@@ 1 Introduction
Finite-time Lyapunov exponents are a quite useful tool to detect local stability/instability of a stochastic system. Negative finite-time Lyapunov exponents indicate attraction whereas positive ones indicate that nearby solutions tend to separate on a finite time horizon. These have been investigated for SDEs [AK:84, CDLR17, DELR:18, BBBE:25] and SPDEs with cubic nonlinearities such as Allen-Cahn or Swift-Hohenberg in [BlEnNe:21, BlNe:23, BB:25, B:25]. Quadratic nonlinearities are significantly more difficult, as cubic nonlinearities often add additional stability and disspation to the system. To our best knowledge, this is the first work that investigates finite-time Lyapunov exponents (FTLE) for SPDEs with quadratic nonlinearities, such as the stochastic Burgers equation.
@@
Based on ideas of our previous work [BlNe:23] for cubic nonlinearites we restrict ourselves to equations close to a change of stability, where the theory of amplitude equations (AE) allows to reduce the dynamics to an SDE on the dominating modes that change stability. This approximation is a well established tool and many results were published on this topic such as [BlHa:04, Bl:07, BlHaPa:15, BlMo:09, BFu:20]. Most of these results use an approximation over a long but finite time-scale, which is the natural slow time-scale on which the interesting dominating dynamics evolve. There are also results available for approximation of invariant measures via AE, see for example [BlHa:04]. Nevertheless, an approximation of Lyapunov-exponents, which takes into account the limit of time to infinity, seems to be out of reach by this method. Therefore we focus on FTLE, since a change of sign in the FTLEs detects a transition which is not the case for asymptotic Lyapunov exponents [CDLR17].
@@
In an informal way our main theorem 4.6 approximates the FTLEs for the full SPDE via the FTLE of the AE. In contrast to our previous work [BlNe:23] we present a fairly general approximation result in Theorem 4.4 that reduces the error term between the finite-time Lyapunov exponents of the SPDE and the amplitude SDE to questions of stability. We believe that this abstract result can be very useful for numerical purposes. We plan to address this aspect in a future work for the stochastic Burgers equation exploiting the techniques in [BlJ:13]. In this work we need precise error analysis of the approximation for amplitude equations (see Theorem 3.9) and the corresponding linearizations (see Theorem 4.2). This allows us to state implications on the bifurcation (positive or negative finite time Lyapunov exponent) once we know the FTLE for the AE. While the approximation result is fairly general, the true bottleneck for a description of bifurcation of FTLE are the lack of results for AE. Let us remark that for 1D SDEs, there are many results for FTLE especially with additive noise [CDLR17, B:25]. The key point is that the solution of the linearized equation is usually given as an explicit exponential, which allows for direct computations. For FTLEs in 2D, we plan to use the techniques in [DELR:18, BBBE:25] and also to exploit symmetries which might help to further reduce the dimension.
@@
In our examples, we mainly focus on a Burgers-type equation with a one-dimensional kernel, but this is mainly due to the availability of results on FTLE for 1D SDEs. We can also treat equations like Navier-Stokes, although there is no bifurcation, and no reduction to a dominant space $\mathcal{N}$. Here we could focus on the onset of convection in Rayleigh-Bénard, where Navier-Stokes is coupled to a heat equation. The operator in that equation is not self-adjoint and we have to modify our approach, but it does not seem to be out of reach, see Subsection 6.3 for a brief discussion. Also models like Kuramoto-Sivashinsky or KPZ-type equations would fall into the scope of our result.
The combination of cubic and quadratic nonlinearities in the model are also possible. Moreover higher order terms in the nonlinearity of the SPDE would do not affect the approximation via AE too much. In order to focus on the main features and challenges posed by the quadratic nonlinearity, and not get overwhelmed by technical details we treat only quadratic nonlinearities here in this paper.
@@ 2 Setting and Assumptions
$\begin{cases}{\textnormal{d}}u=[Au+\nu u+B(u,u)]\penalty 10000\ {\textnormal{d}}t+\sigma{\textnormal{d}}W_{t}\\ u(0)=u_{0}\in X.\end{cases}$ (2.1)
@@
For the $\mathcal{O}$-notation here we use that an $X$-valued process $M$ is $\mathcal{O}(f)$ for a term $f$ on a possibly random interval $I$, if for all probabilities $p\in(0,1)$ there is a constant $C_{p}>0$ such that $\mathbb{P}(\sup_{t\in I}\|M(t)\|_{X}\leq C_{p}f)\geq p$.
For time independent quantities we use the similar notation without the supremum in time. If the process $M$ and the bound $f$ depends on some small quantity $\varepsilon>0$ we assume that the constant $C_{p}$ is independent of $\varepsilon\in(0,\varepsilon_{0}]$ for some fixed $\varepsilon_{0}>0$.
We also use the abbreviation that a process $X_{\varepsilon}$ is $\mathcal{O}(\varepsilon^{n-})$ if for all $\kappa>0$ we have $X_{\varepsilon}=\mathcal{O}(\varepsilon^{n-\kappa})$.
@@ 3 Approximation with amplitude equations
A maximal local solution of (2.1) is an $X$-valued continuous process $u$ defined for times up to stopping time $\tau^{\text{ex}}$ satisfying the variation of constants formula for $t\in[0,\tau^{\star})$
$u(t)=e^{tA}u_{0}+\nu\int_{0}^{t}e^{A(t-s)}u(s)\penalty 10000\ {\textnormal{d}}s+\int_{0}^{t}e^{A(t-s)}B(u(s),u(s))\penalty 10000\ {\textnormal{d}}s+\sigma Z(t),\penalty 10000\ \penalty 10000\ t>0.$ (3.1)
@@
(Approximation) For the approximation of the SPDE (2.1) with a suitable SDE we make the following assumptions:
@@
$\leq C\Big(\sum\limits_{k=1}^{\infty}\|Q^{1/2}_{c}e_{k}\|^{2}_{X}\Big)^{p/2}\mathbb{E}\sup\limits_{T\in[0,\tau^{\star}]}\|A^{-1}_{s}U_{s}(T)\|^{p}_{X}$
@@ 4.1 Linearization
    The term $\varepsilon^{-1}$ appearing in front of the nonlinearity in (4.1) is large and has to be handled by completely different tools than the ones in [BlNe:23] where this factor canceled to the cubic structure of the nonlinearity.
@@
$\|V_{s}(T)\|\leq C_{p}\varepsilon^{1-\kappa},\quad\|V_{c}(T)-\varphi(T)\|\leq C_{p}\varepsilon^{1-\kappa}\quad\text{ for }T\in[T_{\varepsilon},T],$
@@
We further use that $U_{c}=\mathcal{O}(1)$, and $U_{s}=\mathcal{O}(\varepsilon^{0-})$ on the slow time scale $0,T_{0}]$ from Theorem 3.9 and the definition of the stopping time $\tau^{v}$ to bound $V_{s}$. Therefore we infer on $0,\tau^{v}]$ that
@@
$+\varepsilon^{2\alpha}\int_{0}^{T}e^{-\mu(T-S)\varepsilon^{-2}}(T-S)^{-\alpha}[\mathcal{O}(\varepsilon^{-1-}(r_{c}+r_{s}))]\penalty 10000\ {\textnormal{d}}S$
$\leq e^{-\mu T\varepsilon^{-2}}\|V_{s}(0)\|+\mathcal{O}(\varepsilon^{1-})(r_{c}+r_{s}).$ (4.4)
@@
$\sup_{[0,\tau^{v}]}\|V_{s}(T)\|\leq\frac{1}{2}r_{s}\quad\text{and}\quad\sup_{[T_{\varepsilon},\tau^{v}]}\|V_{s}(T)\|=\mathcal{O}(\varepsilon^{1-})$
@@
$\int_{0}^{T}B_{c}(U_{s},V_{s})\penalty 10000\ {\textnormal{d}}S=\int_{0}^{T}B_{c}(\mathcal{O}(\varepsilon^{0-}),e^{-\mu T\varepsilon^{-2}}\mathcal{O}(1)+\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c}))\penalty 10000\ {\textnormal{d}}S=\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c}).$
Similarly,
$\int_{0}^{T}B_{c}(U_{c},V_{s})\penalty 10000\ {\textnormal{d}}S=\int_{0}^{T}B_{c}(\mathcal{O}(1),e^{-\mu T\varepsilon^{-2}}\mathcal{O}(1)+\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c}))\penalty 10000\ {\textnormal{d}}S=\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c})$
and using Corollary 3.10 and $\sigma\leq\varepsilon^{2}$
@@
Thus from (4.5) we have on $0,\tau^{v}]$
$\|V_{c}(T)\|\leq\mathcal{O}(1)\int_{0}^{T}\|V_{c}(S)\|\penalty 10000\ {\textnormal{d}}S+\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c})$
Using Gronwall’s inequality and $\int_{0}^{T_{0}}\|\tilde{Z}_{\varepsilon}(T)\|{\textnormal{d}}S=\mathcal{O}(\varepsilon)$ we have
$\sup_{[0,\tau_{v}]}\|V_{c}(T)\|=\mathcal{O}(1)\|V_{c}(0)\|+\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c})<r_{c}$
@@
$V_{s}=\mathcal{O}(\varepsilon^{1-})\text{ on }[T_{\varepsilon},T_{0}]\text{ and }V_{c}=\mathcal{O}(1)\text{ on }[0,T_{0}].$
@@ 4.2 Approximation of finite-time Lyapunov exponents
    $\sup\limits_{[T_{\varepsilon},T]}\|V_{s}(T)\|=\mathcal{O}(\varepsilon^{1-}).$
    This means that provided the initial conditions are $u(0)=\varepsilon a_{0}+\mathcal{O}(\varepsilon^{2})$ we have $K_{\mathcal{N}}=\mathcal{O}(\varepsilon^{1-})$ but for $K_{X}$ since $\|V_{s}(T)\|=\mathcal{O}(\varepsilon^{1-})$ for $T\in[T_{\varepsilon},T_{0}]$ this is not the case. To be more precise, using (4.4) we have for an arbitrarily small $\kappa>0$
@@ 4.3 Main result
Fix $T_{0}>0$, $\alpha\in(0,1)$, and small $\kappa>0$. Moreover, consider initial conditions $a_{0}=a_{0}(\omega)=\mathcal{O}(1)$ for the AE (3.11) and $u_{0}=\varepsilon a_{0}+\varepsilon^{2}\psi_{0}$ with $\psi_{0}\in\mathcal{S}$ an $\psi=\mathcal{O}(1)$ for the SPDE (2.1). For any probability $p\in(0,1)$ there is a set $\Omega_{p}$ with $\mathbb{P}(\Omega_{p})>p$, and $\varepsilon_{0}>0$ and a constant $C>0$ such that for all $T\in[\varepsilon^{\alpha},T_{0}]$
@@
5 Bifurcations for one-dimensional kernel

### v2

@@ Abstract
We analyze stochastic partial differential equations (SPDEs) with quadratic nonlinearities close to a change of stability / phase-transition. To this aim we compute finite-time Lyapunov exponents (FTLEs), observing a change of sign based on the interplay between the distance towards the bifurcation and the noise intensity. A technical challenge is to provide a suitable control of the nonlinear terms coupling the dominant and stable modes of the SPDE and of the corresponding linearization. In order to illustrate our results we apply them to the stochastic Burgers equation.
Keywords: phase transitions, finite-time Lyapunov exponents, Burgers equation.
@@ 1 Introduction
Finite-time Lyapunov exponents are a very useful tool to detect local stability/instability of a stochastic system. Negative finite-time Lyapunov exponents indicate attraction whereas positive ones indicate that nearby solutions tend to separate on a finite time horizon. These have been investigated for SDEs [AK:84, CDLR17, DELR:18, BBBE:25] and SPDEs with cubic nonlinearities such as Allen-Cahn or Swift-Hohenberg in [BlEnNe:21, BlNe:23, BB:25, B:25]. Quadratic nonlinearities are significantly more difficult, as cubic nonlinearities often add additional stability and dissipation to the system. To our best knowledge, this is the first work that investigates finite-time Lyapunov exponents (FTLE) for SPDEs with quadratic nonlinearities, such as the stochastic Burgers equation where the tools developed in [BlEnNe:21, BlNe:23, BB:25] break down. The quadratic nonlinearity introduces a coupling between the dominant and stable modes of the SPDE and its linearization which does not occur for stable cubic nonlinearities. Therefore removing the stable modes using averaging techniques and getting closed equations for the dominant modes is challenging and requires a suitable Itô trick together with a careful analysis of the coupling.
@@
In order to study a bifurcation via Lyapunov exponents, we consider equations close to a change of stability, where the theory of amplitude equations (AE) allows to reduce the dynamics to an SDE on the dominating modes. This approximation is a well established tool and many results were published on this topic such as [BlHa:04, Bl:07, BlHaPa:15, BlMo:09, BFu:20]. Most of these results use an approximation over a long but finite time-scale, which is the natural slow time-scale on which the interesting dominating dynamics evolve. There are also results available for approximation of invariant measures via AE, see for example [BlHa:04]. Nevertheless, an approximation of Lyapunov-exponents, which takes into account the limit of time to infinity, seems to be out of reach at the moment. Therefore we focus on FTLE, since a change of sign in the FTLEs detects a transition which is not the case for asymptotic Lyapunov exponents [CDLR17].
@@
In an informal way our main result, Theorem 4.6 approximates the FTLEs for the full SPDE via the FTLE of the AE. In contrast to our previous work [BlNe:23] we present a fairly general approximation result in Theorem 4.4 that reduces the error term between the finite-time Lyapunov exponents of the SPDE and the amplitude SDE to questions of stability. We believe that this abstract result can be very useful for numerical purposes. We plan to address this aspect in a future work for the stochastic Burgers equation exploiting the techniques in [BlJ:13]. In this work we need precise error analysis of the approximation for amplitude equations (see Theorem 3.9) and the corresponding linearizations (see Theorem 4.2). This allows us to state implications on the bifurcation (positive or negative finite time Lyapunov exponent) once we know the FTLE for the AE. While our main approximation result is fairly general, we also need results on FTLEs for AE. Let us remark that for 1D SDEs, there are many results for FTLE especially with additive noise [CDLR17, B:25]. The key point is that the solution of the linearized equation is usually given as an explicit exponential, which allows for direct computations. For FTLEs in 2D, we plan to use the techniques in [DELR:18, BBBE:25] and also to exploit symmetries which might help to further reduce the dimension.
@@
In our examples, we mainly focus on a Burgers-type equation with a one-dimensional kernel, but this is mainly due to the availability of results on FTLE for 1D SDEs. We can also treat equations like Navier-Stokes, although there is no bifurcation, and no reduction to a dominant space $\mathcal{N}$. Here we could focus on the onset of convection in Rayleigh-Bénard, which consists of a Navier-Stokes equation coupled to a heat equation. The operator in this model is not self-adjoint and therefore we have to modify our approach, but this does not seem to be out of reach and will be treated in a future work, see Subsection 6.3 for a brief discussion. Moreover, models like Kuramoto-Sivashinsky or KPZ-type equations would fall into the scope of our result. Another example is a surface growth model in [BlMo:09] with periodic boundary conditions where the kernel is two-dimensional, here in some cases we might exploit symmetries like polar coordinates to reduce the AE from 2D to 1D. The combination of cubic and quadratic nonlinearities in the model are also possible. Moreover higher order terms in the nonlinearity of the SPDE would not affect the approximation via AE too much. In order to focus on the main features and challenges posed by the quadratic nonlinearity, and not get overwhelmed by technical details we treat only quadratic nonlinearities here in this paper.
@@ Acknowledgments
Conflict of interest statement
The authors have no conflicts of interest to declare. All co-authors have seen and agree with the contents of the manuscript and there is no financial interest to report.
Data availability
No data was used for the research described in the article.
Ethics declaration
Not applicable.
@@ 2 Setting and Assumptions
$\begin{cases}{\textnormal{d}}u=[Au+\nu u+B(u,u)]\penalty 10000\ {\textnormal{d}}t+\sigma{\textnormal{d}}W(t)\\ u(0)=u_{0}\in X.\end{cases}$ (2.1)
@@
For the $\mathcal{O}$-notation here we use that an $X$-valued process $M$ is $\mathcal{O}(f)$ for a term $f$ on a possibly random interval $I$, if for all probabilities $p\in(0,1)$ there is a constant $C_{p}>0$ such that $\mathbb{P}(\sup_{t\in I}\|M(t)\|_{X}\leq C_{p}f)\geq p$. For time independent quantities we use the similar notation without the supremum in time. If the process $M$ and the bound $f$ depends on some small quantity $\varepsilon>0$ we assume that the constant $C_{p}$ is independent of $\varepsilon\in(0,\varepsilon_{0}]$ for some fixed $\varepsilon_{0}>0$. We also use the abbreviation that a process $M_{\varepsilon}$ is $\mathcal{O}(\varepsilon^{n-})$ if for all $\kappa>0$ we have $M_{\varepsilon}=\mathcal{O}(\varepsilon^{n-\kappa})$.
@@ 3 Approximation with amplitude equations
A maximal local solution of (2.1) is an $X$-valued continuous process $u$ defined for times up to stopping time $\tau^{\text{ex}}$ satisfying the variation of constants formula for $t\in[0,\tau^{\text{ex}})$
$u(t)=e^{tA}u_{0}+\nu\int_{0}^{t}e^{A(t-s)}u(s)\penalty 10000\ {\textnormal{d}}s+\int_{0}^{t}e^{A(t-s)}B(u(s),u(s))\penalty 10000\ {\textnormal{d}}s+\sigma Z(t),\penalty 10000\ \penalty 10000\ t>0,$ (3.1)
@@
(Approximation) We make the following assumptions:
@@
$\leq C\Big(\sum\limits_{k=1}^{\infty}\|Q^{1/2}_{c}e_{k}\|^{2}_{X}\Big)^{p/2}\mathbb{E}\sup\limits_{T\in[0,\tau^{\star}]}\|A^{-1}_{s}U_{s}(T)\|^{p}_{X}.$
@@ 4.1 Linearization
    The term $\varepsilon^{-1}$ appearing in front of the nonlinearity in (4.1) is large and has to be handled by completely different tools than the ones in [BlNe:23] where this factor canceled due to the cubic structure of the nonlinearity.
@@
$\|V_{s}(T)\|\leq C_{p}\varepsilon,\quad\|V_{c}(T)-\varphi(T)\|\leq C_{p}\varepsilon^{1-\kappa}\quad\text{ for }T\in[T_{\varepsilon},T],$
@@
We further use that $U_{c}=\mathcal{O}(1)$, and $U_{s}=\mathcal{O}(\varepsilon^{0-})$ on the slow time scale $0,T_{0}]$ from Theorem 3.9 and the definition of the stopping time $\tau^{v}$ to obtain
$\|B_{s}(U,V)\|_{Y}\leq\mathcal{O}(1)(r_{c}+r_{s})+\mathcal{O}(\varepsilon^{1-})(r_{c}+r_{s}).$
Therefore we infer on $0,\tau^{v}]$ that
@@
$+\varepsilon^{2\alpha}\int_{0}^{T}e^{-\mu(T-S)\varepsilon^{-2}}(T-S)^{-\alpha}[\mathcal{O}(\varepsilon^{-1}(r_{c}+r_{s}))]\penalty 10000\ {\textnormal{d}}S$
$\leq e^{-\mu T\varepsilon^{-2}}\|V_{s}(0)\|+\mathcal{O}(\varepsilon)(r_{c}+r_{s}).$ (4.4)
@@
$\sup_{[0,\tau^{v}]}\|V_{s}(T)\|\leq\frac{1}{2}r_{s}\quad\text{and}\quad\sup_{[T_{\varepsilon},\tau^{v}]}\|V_{s}(T)\|=\mathcal{O}(\varepsilon(r_{c}+r_{s}))$
@@
Moreover, we can modify (4.4) to obtain
$\|V_{s}(T)\|\leq e^{-\mu T\varepsilon^{-2}}\|V_{s}(0)\|+\mathcal{O}(\varepsilon^{2\alpha-1})\int_{0}^{T}e^{-\mu(T-S)\varepsilon^{-2}}(T-S)^{-\alpha}\|V_{c}(S)\|\penalty 10000\ {\textnormal{d}}S+\mathcal{O}(\varepsilon)r_{s}.$ (4.5)
@@
$\int_{0}^{T}B_{c}(U_{s},V_{s})\penalty 10000\ {\textnormal{d}}S=\int_{0}^{T}B_{c}(\mathcal{O}(\varepsilon^{0-}),e^{-\mu T\varepsilon^{-2}}\mathcal{O}(1)+\mathcal{O}(\varepsilon^{1-})(r_{s}+r_{c}))\penalty 10000\ {\textnormal{d}}S=\mathcal{O}(\varepsilon)(r_{s}+r_{c}).$
Similarly, using additionally (4.5) we obtain
$\varepsilon^{-1}\int_{0}^{T}B_{c}(U_{c},V_{s})\penalty 10000\ {\textnormal{d}}S=$ $\mathcal{O}(\varepsilon^{-1})\int_{0}^{T}e^{-\mu S\varepsilon^{-2}}\penalty 10000\ {\textnormal{d}}S+\mathcal{O}(1)r_{s}$
$+\mathcal{O}(\varepsilon^{2\alpha-2})\int_{0}^{T}\int_{0}^{S}e^{-\mu(S-s)\varepsilon^{-2}}(S-s)^{-\alpha}\|V_{c}(s)\|\penalty 10000\ {\textnormal{d}}s\penalty 10000\ {\textnormal{d}}S$
Using Fubini we obtain
$\varepsilon^{-1}\int_{0}^{T}B_{c}(U_{c},V_{s})\penalty 10000\ {\textnormal{d}}S=\mathcal{O}(1)\int_{0}^{T}\|V_{c}(S)\|{\textnormal{d}}S+\mathcal{O}(1)r_{s}.$
Finally, using Corollary 3.10 and $\sigma\leq\varepsilon^{2}$
@@
Thus from (4.6) we have on $0,\tau^{v}]$ for small $\varepsilon$
$\|V_{c}(T)\|\leq\mathcal{O}(1)(1+\varepsilon^{-1}\|\tilde{Z}_{\varepsilon}\|)\int_{0}^{T}\|V_{c}(S)\|\penalty 10000\ {\textnormal{d}}S+\mathcal{O}(1)r_{s}+\mathcal{O}(\varepsilon)r_{c}.$
Using Gronwall’s inequality and $\int_{0}^{T_{0}}\|\tilde{Z}_{\varepsilon}(S)\|{\textnormal{d}}S=\mathcal{O}(\varepsilon)$ we have
$\sup_{[0,\tau_{v}]}\|V_{c}(T)\|=\mathcal{O}(1)\|V_{c}(0)\|+\mathcal{O}(1)r_{s}+\mathcal{O}(\varepsilon)r_{c}<r_{c}$
@@
$V_{s}=\mathcal{O}(\varepsilon)\text{ on }[T_{\varepsilon},T_{0}]\text{ and }V_{c}=\mathcal{O}(1)\text{ on }[0,T_{0}].$
@@ 4.2 Approximation of finite-time Lyapunov exponents
    $\sup\limits_{[T_{\varepsilon},T]}\|V_{s}(T)\|=\mathcal{O}(\varepsilon).$
    This means that provided the initial conditions are $u(0)=\varepsilon a_{0}+\mathcal{O}(\varepsilon^{2})$ we have $K_{\mathcal{N}}=\mathcal{O}(\varepsilon^{1-})$ but for $K_{X}$ since $\|V_{s}(T)\|=\mathcal{O}(\varepsilon)$ for $T\in[T_{\varepsilon},T_{0}]$ this is not the case. To be more precise, using (4.4) we have for an arbitrarily small $\kappa>0$
@@ 4.3 Main result
Fix $T_{0}>0$, $\alpha\in(0,1)$, and small $\kappa>0$. Moreover, consider initial conditions $a_{0}=a_{0}(\omega)=\mathcal{O}(1)$ for the AE (3.11) and $u_{0}=\varepsilon a_{0}+\varepsilon^{2}\psi_{0}$ with $\psi_{0}\in\mathcal{S}$ and $\psi=\mathcal{O}(1)$ for the SPDE (2.1). For any probability $p\in(0,1)$ there is a set $\Omega_{p}$ with $\mathbb{P}(\Omega_{p})>p$, and $\varepsilon_{0}>0$ and a constant $C>0$ such that for all $T\in[\varepsilon^{\alpha},T_{0}]$
@@
5 Example: Bifurcations for one-dimensional kernel